OpenStudy (anonymous):

factor the polynomial completely 2u^2+ 2v^2-18

Parth (parthkohli):

Factor out the \(2\) to start with.

OpenStudy (anonymous):

2(U^2 V^2-18) then what

OpenStudy (anonymous):

i need help

OpenStudy (mathmate):

2(U^2+V^2-9)

OpenStudy (anonymous):

\[ 2\left(u^2+v^2 -9 \right) \]

OpenStudy (anonymous):

i goty that

OpenStudy (mathmate):

If you push it, you can factor (u^2-9) as (u+3)(u-3)

OpenStudy (anonymous):

but wait

OpenStudy (anonymous):

a 2(u^2 +V^2 -9) b 2uv(u+v-9) C 2(U^2+V^20-18) D 2UV(u+v)-18 these are the answer choices

OpenStudy (anonymous):

is it b by anychance because if you use foil you get the same problem you start with

OpenStudy (mathmate):

didn't elissaab give you the exact answer which corresponds to one of the choices?

OpenStudy (anonymous):

but thats not factoring compelety

OpenStudy (anonymous):

is that factoring completely confused

OpenStudy (anonymous):

elisaab"s answer is correct

OpenStudy (anonymous):

its A

OpenStudy (mathmate):

The best you can do, further, is \[2((u+3)(u-3)+v^2)\] which is not a complete factorization either, and is not one of the choices.

OpenStudy (mathmate):

Factoring completely does not imply that all factors have to be linear (degree 1).

OpenStudy (anonymous):

telltoamit Hi its Tahmina again

OpenStudy (mathmate):

It just means that each factor cannot be further factored.

OpenStudy (anonymous):

I mean the further in education u are in the more you know

OpenStudy (mathmate):

Either she didn't tell you, or you didn't hear. (This statement is always true, or a tautology).

OpenStudy (anonymous):

i didnt know about i jst did that just right now

OpenStudy (anonymous):

so its A rite because i had that and i wasnt sure

OpenStudy (anonymous):

@Tahmina : Hi! How u doin? :)

OpenStudy (mathmate):

Another example: \[x^3-y^3=(x-y)(x^2+xy+y^2)\] is factored completely.

OpenStudy (anonymous):

Yes, its A

OpenStudy (anonymous):

so solve the equation below 49N^2=100

OpenStudy (anonymous):

how far hv u gone in it?

OpenStudy (anonymous):

is it 10/7 , -10/7

OpenStudy (anonymous):

awsumm, correct

OpenStudy (anonymous):

yaa

OpenStudy (anonymous):

you know school is starting and I want to get everything correct if i dont and get a bad grade then Im kicked out of ARP

OpenStudy (anonymous):

something like thAT

OpenStudy (anonymous):

wuzz ARP?

OpenStudy (anonymous):

Advanced Regents Placement i was in it for 2 years already

OpenStudy (anonymous):

and if im not in ARP I DONT NO HOW I WILL BE arista

OpenStudy (anonymous):

okk

OpenStudy (anonymous):

i WANT TO BE IN ARISTA FOR 3 YEARS I WAS ALREADY FOR 2 YEARS

OpenStudy (anonymous):

OKAY BYE :)

OpenStudy (anonymous):

bbyezz

OpenStudy (anonymous):

this one?